OpenStudy (anonymous):

Multiply x-1/4x+4 . x^2-2x-3/x^2+x-2 simplify your answer as much as you can

OpenStudy (wolfe8):

\[\frac{ x-1 }{ 4x+4 }\times \frac{ x ^{2} -2x-3}{ x ^{2}+x-2 }\] What you do is just multiply the numerators together ontop and denominators together below. Do you get that?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

-7 on top

OpenStudy (anonymous):

correct?

OpenStudy (wolfe8):

How did you get a constant?

OpenStudy (anonymous):

multiply the numbers on top

OpenStudy (anonymous):

i got it

OpenStudy (anonymous):

THanks

OpenStudy (wolfe8):

\[\frac{ (x-1)(x ^{2}-2x-3) }{ (4x+4)(x ^{2}+x-2) }\]